# Wi-Fi Radar via Over-the-Air Referencing: Bridging Wi-Fi Sensing and Bistatic Radar

### v2

@@ VI Delay–Doppler Response Estimation of Moving Human Target
This section investigates delay–Doppler response estimation of a moving human target using phase-coherent Wi-Fi CIRs. To this end, we first characterize the impact of irregular temporal sampling in practical Wi-Fi systems and introduce time-domain resampling. We then construct a delay–Doppler representation and demonstrate range–Doppler estimation through experiments.
@@ VI-A Inter-Packet Timing Variations in Contention-Based Wi-Fi
Although frame transmissions are nominally scheduled at regular intervals, the actual observation times at the Rx are not guaranteed to be uniformly spaced. This is due to contention-based medium access, occasional missed receptions, and the removal of outliers.
Figure 15: Temporal variation and empirical distribution of the observed inter-packet intervals. Although the nominal sampling interval is $1\text{\,}\mathrm{ms}$, the intervals exhibit noticeable fluctuations, with a median of $1.069\text{\,}\mathrm{ms}$.
This timing variability was experimentally examined in Fig. 15. Fig. 15LABEL:sub@fig:intersample_time illustrates the inter-packet intervals computed from the observed measurements in the experiment described in Section V-D, where packet transmissions were intended to follow a $1\text{\,}\mathrm{ms}$ interval. The inter-packet intervals were spread over a range from approximately $0.5\text{\,}\mathrm{ms}$ to $6.5\text{\,}\mathrm{ms}$, indicating significant deviations from the nominal sampling period.
Fig. 15LABEL:sub@fig:intersample_time_cdf shows the corresponding empirical cumulative distribution function (CDF) of the inter-packet intervals. The observed intervals exhibited a median of $1.069\text{\,}\mathrm{ms}$, with a median absolute deviation of $0.017\text{\,}\mathrm{ms}$, indicating small but non-negligible timing jitter around the nominal sampling interval.
@@ VI-B Uniform Temporal Resampling for Phase-Coherent Doppler Estimation
Since Doppler estimation generally requires uniformly spaced observations in time, a uniform time grid is constructed, where the grid spacing $\varDelta t$ is set to the median inter-packet interval, i.e., $\varDelta t=$1.069\text{\,}\mathrm{ms}$\$ in the present measurements. Each observed CIR snapshot is first associated with its temporally nearest grid point. The remaining unassociated grid points are then filled by interpolation: the magnitude is interpolated linearly, while the phase is interpolated after phase unwrapping so as to preserve temporal continuity. Although more advanced treatments of non-uniform sampling are possible, the above procedure is sufficient for accurate phase-coherent Doppler estimation in this work. Although more advanced treatments of non-uniform sampling are possible, the above grid-based resampling procedure is sufficient for accurate Doppler estimation in this work.
@@ VI-C Delay–Doppler Representation and Doppler Interpretation
The delay–Doppler representation is obtained by Fourier transforming the CIR $h(\tau,t)$ with respect to time $t$, as is well established in the literature [28, 27]. While Section V-D illustrated Doppler behavior over a short time interval at a fixed delay, the delay–Doppler representation generalizes this observation by characterizing Doppler frequency as a function of delay. In practice, we apply an STFT along the time axis to the phase-coherent and uniformly resampled CIR $h(\tau,t)$, yielding a time-varying delay–Doppler representation $s(\tau,\nu;t)$, where $t$ denotes the center time of the analysis window. [^2] The CIR $h(\tau,t)$ may be taken either before or after clutter removal in Section IV-C, since clutter components are suppressed on a per-STFT-segment basis. This representation jointly captures the temporal evolution of multipath delays and their associated Doppler frequencies.
Although the experiment employs a bistatic configuration, the Tx–Rx separation is sufficiently small that the geometry can be approximated as monostatic. In this limit, the bistatic range $d_{\text{Tx--target--Rx}}$ reduces to twice the one-way Tx/Rx–target distance, and the bistatic range rate becomes
$\dot{d}_{\text{Tx--target--Rx}}=2v,$ (13)
where $v$ denotes the radial velocity defined as the time derivative of the one-way Tx/Rx–target distance. Substituting (13) into (12) yields
$\nu=-{2v}/{\lambda}.$ (14)
For ease of interpretation, we therefore employ an effective radial velocity scaled as $\nu\lambda/2$ in the following discussion, instead of the bistatic range rate.
@@ VI-D Experimental Results: Range–Doppler Estimation
The main STFT parameters are summarized in Table III. Similar to the Fourier transform from the frequency domain to the delay domain in (3), a Hann window was applied to each STFT segment of length $256$ samples to suppress spectral leakage, with an overlap of $224$ samples (87.587.5%), corresponding to a hop size of $32\text{\,}\mathrm{ms}$. With an STFT segment of $256$ samples and a time grid of $\varDelta t\approx$1\text{\,}\mathrm{ms}$\$, the resulting Doppler frequency resolution, determined by the effective STFT window duration, is approximately $\varDelta\nu\approx$3.9\text{\,}\mathrm{Hz}$\$, which is sufficient to resolve human motion-induced Doppler components. In addition, Doppler-domain oversampling with a factor of $8$ was performed by zero-padding in the time domain. Furthermore, the mean value of each STFT segment was removed to mitigate DC components. This operation can be interpreted as applying clutter removal described in Section IV-C at the level of individual STFT segments.
@@
A representative example of the range–Doppler estimation ${s}(\tau,\nu;t)$, obtained from the moving human target measurements described in Section V-D, was shown in Fig. 16. The evolution of the delay–Doppler structure over time clearly reflected the target motion relative to the Tx–Rx pair.
@@
Fig. 17 shows the estimated bistatic range $c\tau^{\star}$ and Doppler frequency $\nu^{\star}$ of the moving human target, together with the corresponding geometry-based values derived from the motion-tracking system. Here, $\tau^{\star}$ and $\nu^{\star}$ were obtained by selecting, at each time instant $t$, the maximum of the delay–Doppler representation $s(\tau,\nu;t)$, i.e.,
@@
Overall, the estimated bistatic range and Doppler frequency exhibited trends that are broadly consistent with the geometry-based reference. It should be noted that the tracking system captures the motion of a marker attached to the target’s head, while the Doppler estimates may be influenced by distributed scattering from multiple body parts. Therefore, perfect agreement is not expected.
More specifically, Fig. 17LABEL:sub@fig:range-time-peak corresponds to the delay–time heatmap shown in Fig. 10LABEL:sub@fig:h_tau-t. The bistatic range shown here is given by $c\tau^{\star}$ extracted from the delay–Doppler representation, rather than by directly taking the peak of the residual CIR $\tilde{h}(\tau,t)$.
In Fig. 17LABEL:sub@fig:doppler-time-peak, the estimated Doppler frequency $\nu^{\star}$ is shown together with the corresponding effective radial velocity $\nu^{\star}\lambda/2$, evaluated using (14), on the right axis. For typical Wi-Fi carrier frequencies, human motion gives rise to Doppler frequencies on the order of several tens of hertz [24, 35]. With a uniform time grid of $\varDelta t\approx$1\text{\,}\mathrm{ms}$\$, the resulting effective Doppler bandwidth of $\pm$500\text{\,}\mathrm{Hz}$\$ is therefore sufficient for capturing such motion-induced Doppler components.
@@
The peak of $s(\tau,\nu;t)$, i.e., $(\tau^{\star},\nu^{\star})$, provides intuitive point estimates of the bistatic range and Doppler frequency, but does not capture the full Doppler energy distribution. To complement this point-wise representation, we additionally examined a Doppler–time representation obtained by incoherently integrating the power $\lvert s(\tau,\nu;t)\rvert^{2}$ along the delay axis, as shown in Fig. 18. This representation visualizes the temporal Doppler energy and exhibits a similar trend to that in Fig. 17LABEL:sub@fig:doppler-time-peak; however, the two representations are not directly comparable, as they are derived through different aggregation procedures.
@@ VI-E Comparison with Frequency-Domain Magnitude-Based Wi-Fi Sensing
For comparison, we also present a representative conventional Wi-Fi sensing result that operates purely in the frequency domain and does not rely on the CIR representation. As discussed in Section II-A, many existing approaches perform sensing directly on the magnitude of the CFR, $\lvert H_{\mathrm{raw}}(f,t)\rvert$. In this paper, PCA is used as a representative example of dimensionality reduction, as in [24].
@@
Among the resulting PC scores, the STFT is applied to the PC3 score, which exhibits clear target-induced temporal variations. It should be noted, however, that the PC capturing motion-induced dynamics is device- and environment-dependent [24, 23], and its identification often relies on empirical or heuristic criteria rather than a physical model. The resulting Doppler-related time–frequency representation is shown in Fig. 19.
@@
A key limitation of this magnitude-based frequency-domain approach is that it does not yield a physically meaningful Doppler frequency, because the underlying propagation-induced phase evolution is not preserved, as explained in Section II-A with reference to [26]. As a result, the spectrum in Fig. 19 is symmetric with respect to frequency sign and effectively collapses positive and negative Doppler components, making it impossible to distinguish approaching and receding motions.
@@
By contrast, the proposed LoSRef-based approach operates on the complex-valued residual CIR, whose Doppler spectrum is inherently asymmetric with respect to frequency sign. This asymmetry preserves Doppler sign information and enables direct discrimination between approaching and receding motions, which is fundamentally impossible with magnitude-based processing.
VII Discussion
The experimental validation was performed using the Intel AX210 chipset. While multiple tools for CSI extraction from IEEE 802.11ax devices exist [36, 37], they are all restricted to the Intel AX2xx family, precluding validation on alternative chipsets. Nevertheless, the LoSRef-based framework is chipset-agnostic, as the subsequent processing after phase-coherent CIR construction follows standard bistatic radar signal processing principles, and is therefore applicable to any system that provides a complex-valued CFR and a CIR with sufficient delay resolution.
Recent studies have shown that range–Doppler information can be extracted from CSI using COTS IEEE 802.11ax Wi-Fi devices under monostatic configurations [38]. However, extending such capabilities to bistatic scenarios remains challenging, primarily due to the lack of a reliable reference signal between the Tx and Rx.
VIII Conclusion
@@
This paper presented a LoSRef-based bistatic Wi-Fi radar framework that enables phase-coherent delay–Doppler analysis using only unmodified COTS Wi-Fi devices. LoSRef leverages the Tx–Rx LoS path as a delay and phase reference; its inherent temporal stability enables consistent phase-coherent alignment across successive packet receptions. As a result, the proposed framework eliminates the need for wired references or dedicated reference antennas and achieves phase-coherent bistatic radar operation with a single-antenna, single-port Rx, thereby breaking a key deployment barrier in existing Wi-Fi radar systems.
@@
Through representative human gait and respiration experiments, we confirmed that phase-coherent bistatic radar functionality can be realized within typically deployed, communication-oriented Wi-Fi systems. Specifically, the proposed framework enables physically meaningful delay and Doppler estimates, including gait direction discrimination and respiration-induced sub-wavelength displacement measurements.
@@
The proposed LoSRef-based framework further suggests that Wi-Fi sensing can move beyond application-specific processing. Once phase-coherent representations are restored, radar signal processing techniques can be directly applied to Wi-Fi sensing, pointing to a fruitful research direction for future ISAC system design.
@@ Acknowledgment
The author would like to thank Mr. Genichiro Shiino for providing the experimental data used in this study and for valuable discussions throughout the course of this work. The author also thanks Prof. Katsuyuki Haneda for his insightful comments on an early draft of this paper, and Prof. Hiromitsu Awano for his valuable comments on vital sensing.

### v3

@@ VI Delay–Doppler Response Estimation of Moving Human Target
This section investigates delay–Doppler response estimation of a moving human target using phase-coherent Wi-Fi CIRs. We first characterize the impact of irregular temporal sampling in practical Wi-Fi systems and introduce time-domain resampling. We then construct a delay–Doppler representation and demonstrate range–Doppler estimation through experiments.
@@ VI-A Inter-Packet Timing Variations in Contention-Based Wi-Fi
Although packet transmissions are scheduled at regular intervals, the actual observation times are not guaranteed to be uniformly spaced. This is due to contention-based medium access, missed receptions, and the removal of outliers.
Figure 15: Empirical CDF of the observed inter-packet intervals. Although the nominal sampling interval is $1\text{\,}\mathrm{ms}$, the intervals exhibit noticeable fluctuations, with a median of $1.069\text{\,}\mathrm{ms}$.
Fig. 15 shows the empirical cumulative distribution function (CDF) of the inter-packet intervals obtained in the experiment described in Section V-D, where packet transmissions were intended to follow a $1\text{\,}\mathrm{ms}$ interval. The observed intervals exhibited a median of $1.069\text{\,}\mathrm{ms}$, with a median absolute deviation of $0.017\text{\,}\mathrm{ms}$, indicating small but non-negligible timing jitter around the nominal sampling interval.
@@ VI-B Uniform Temporal Resampling for Phase-Coherent Doppler Estimation
Since Doppler estimation generally requires uniformly spaced observations in time, a uniform time grid is constructed, where the grid spacing $\varDelta t$ is set to the median inter-packet interval, i.e., $\varDelta t=$1.069\text{\,}\mathrm{ms}$\$ in the present measurements. Each observed CIR snapshot is first associated with its temporally nearest grid point. The remaining grid points are then filled by interpolation: the magnitude is interpolated linearly, while the phase is interpolated after phase unwrapping so as to preserve temporal continuity. Although more advanced treatments of non-uniform sampling are possible, the above procedure is sufficient for accurate phase-coherent Doppler estimation in this work.
@@ VI-C Delay–Doppler Representation and Doppler Interpretation
The delay–Doppler representation is obtained by Fourier transforming the CIR $h(\tau,t)$ with respect to time $t$ [28, 27]. While Section V-D illustrated Doppler behavior over a short time interval at a fixed delay, the delay–Doppler representation generalizes this observation by characterizing Doppler frequency as a function of delay. In practice, we apply an STFT along the time axis to the phase-coherent and uniformly resampled CIR $h(\tau,t)$, yielding a time-varying delay–Doppler representation $s(\tau,\nu;t)$, where $t$ denotes the center time of the analysis window. [^2] The CIR $h(\tau,t)$ may be taken either before or after clutter removal, since clutter components are suppressed on a per-STFT-segment basis. This representation jointly captures the temporal evolution of multipath delays and their associated Doppler frequencies.
Although the experiment employs a bistatic configuration, the Tx–Rx separation is small enough that the geometry can be approximated as monostatic. In this limit, $\nu=-2v/\lambda$, where $v$ denotes the target radial velocity. Accordingly, we interpret the Doppler frequency using the effective radial velocity $\nu\lambda/2$.
@@ VI-D Experimental Results: Range–Doppler Estimation
The main STFT parameters are summarized in Table III. Similar to the Fourier transform from the frequency domain to the delay domain in (3), a Hann window was applied to each STFT segment of length $256$ samples to suppress spectral leakage, with an overlap of $224$ samples (87.587.5%), corresponding to a hop size of $32\text{\,}\mathrm{ms}$. With an STFT segment of $256$ samples and a time grid of $\varDelta t\approx$1\text{\,}\mathrm{ms}$\$, the resulting Doppler frequency resolution, determined by the effective STFT window duration, is approximately $\varDelta\nu\approx$3.9\text{\,}\mathrm{Hz}$\$, which is sufficient to resolve human motion-induced Doppler components. In addition, Doppler-domain oversampling with a factor of $8$ was performed by zero-padding in the time domain. Furthermore, the mean value of each STFT segment was removed to mitigate DC components. This operation corresponds to applying clutter removal described in Section IV-C at the level of individual STFT segments.
@@
A representative example of the range–Doppler estimation ${s}(\tau,\nu;t)$, obtained from the moving human target measurements described in Section V-D, was shown in Fig. 16. The evolution of the delay–Doppler structure over time reflected the target motion relative to the Tx–Rx pair. Fig. 17 shows the estimated bistatic range $c\tau^{\star}$ and Doppler frequency $\nu^{\star}$ of the moving human target, together with the corresponding geometry-based values derived from the motion-tracking system. Here, $\tau^{\star}$ and $\nu^{\star}$ were obtained by selecting, at each time instant $t$, the maximum of the delay–Doppler representation $s(\tau,\nu;t)$, i.e.,
@@
Overall, the estimated bistatic range and Doppler frequency are broadly consistent with the geometry-based reference. It should be noted that the tracking system captures the motion of a marker attached to the target’s head, while the Doppler estimates may be influenced by distributed scattering from multiple body parts. Therefore, perfect agreement is not expected. More specifically, Fig. 17LABEL:sub@fig:range-time-peak corresponds to the delay–time heatmap in Fig. 10LABEL:sub@fig:h_tau-t. The bistatic range shown here is given by $c\tau^{\star}$ extracted from the delay–Doppler representation, rather than the peak of the residual CIR $\tilde{h}(\tau,t)$.
In Fig. 17LABEL:sub@fig:doppler-time-peak, the estimated Doppler frequency $\nu^{\star}$ is shown together with the effective radial velocity $\nu^{\star}\lambda/2$. For typical Wi-Fi carrier frequencies, human motion gives rise to Doppler frequencies on the order of several tens of hertz [24, 35]. With a uniform time grid of $\varDelta t\approx$1\text{\,}\mathrm{ms}$\$, the resulting effective Doppler bandwidth of $\pm$500\text{\,}\mathrm{Hz}$\$ is therefore sufficient to capture such motion-induced Doppler components.
@@
The peak of $s(\tau,\nu;t)$, i.e., $(\tau^{\star},\nu^{\star})$, provides point estimates of the bistatic range and Doppler frequency, but does not capture the full Doppler energy distribution. To complement this point-wise representation, we additionally examined a Doppler–time representation obtained by incoherently integrating the power $\lvert s(\tau,\nu;t)\rvert^{2}$ along the delay axis, as shown in Fig. 18. This representation visualizes the temporal Doppler energy and exhibits a similar trend to that in Fig. 17LABEL:sub@fig:doppler-time-peak; however, the two representations are not directly comparable, as they are derived through different aggregation procedures.
@@ VI-E Comparison with Frequency-Domain Magnitude-Based Wi-Fi Sensing
For comparison, we present a conventional Wi-Fi sensing result that operates in the frequency domain and does not rely on the CIR representation. As discussed in Section II-A, many existing approaches perform sensing directly on the magnitude of the CFR, $\lvert H_{\mathrm{raw}}(f,t)\rvert$. In this paper, PCA is used as a representative example of dimensionality reduction, as in [24].
@@
Among the resulting PC scores, the STFT is applied to the PC3 score, which exhibits target-induced temporal variations. It should be noted, however, that the PC capturing motion-induced dynamics is device- and environment-dependent [24, 23], and its identification relies on empirical or heuristic criteria rather than a physical model. The resulting Doppler-related time–frequency representation is shown in Fig. 19.
@@
A key limitation of this magnitude-based frequency-domain approach is that it does not yield a physically meaningful Doppler frequency, because the propagation-induced phase evolution is not preserved, as explained in Section II-A with reference to [26]. As a result, the spectrum in Fig. 19 is symmetric with respect to frequency sign and collapses positive and negative Doppler components, making it impossible to distinguish approaching and receding motions.
@@
By contrast, the proposed LoSRef-based approach operates on the complex-valued residual CIR, whose Doppler spectrum is asymmetric with respect to frequency sign. This asymmetry preserves Doppler sign information and enables discrimination between approaching and receding motions, which is impossible with magnitude-based processing.
VII Conclusion
@@
This paper presented a LoSRef-based bistatic Wi-Fi radar framework that enables phase-coherent delay–Doppler analysis using only unmodified COTS Wi-Fi devices. LoSRef leverages the Tx–Rx LoS path as a delay and phase reference; its temporal stability enables phase-coherent alignment across successive packet receptions. As a result, the proposed framework eliminates the need for wired references or dedicated reference antennas and achieves phase-coherent bistatic radar operation with a single-antenna, single-port Rx, thereby breaking a key deployment barrier in existing Wi-Fi radar systems.
@@
Through human gait and respiration experiments, we confirmed that phase-coherent bistatic radar functionality can be realized within typically deployed, communication-oriented Wi-Fi systems. Specifically, the proposed framework enables physically meaningful delay and Doppler estimates, including gait direction discrimination and respiration-induced sub-wavelength displacement measurements.
@@
The LoSRef-based framework further suggests that Wi-Fi sensing can move beyond application-specific processing. Once phase-coherent representations are restored, radar signal processing techniques can be applied to Wi-Fi sensing, pointing to research direction for future ISAC system design.
@@ Acknowledgment
The author would like to thank Mr. Genichiro Shiino for providing the experimental data used in this study and for valuable discussions. The author also thanks Prof. Katsuyuki Haneda for his insightful comments on an early draft of this paper, and Prof. Hiromitsu Awano for his valuable comments on vital sensing experiments. The author used generative AI tools for language editing during the preparation of this manuscript. The author reviewed and takes full responsibility for the final content.